• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name sarty

    Log entry time 16:02:08 on May20,2008

    Entry number 238970

    keyword=Shift Summary

    Day Shift Summary

    Shift Workers: A. Sarty, H. Yao (A. Saha - gone home: ill)

    0800: MCC no beam - lost orbit locks

    Change HRS to be ready for: Kin2, delta = +2%
    (LHRS: p = 0.5535 MeV/c) with 6 cm LH2 target in place.

    Quick thunderstorm rolls through...takes entire machine down (before beam could be restored). Begin a Controlled Access during machine recovery time.

    14:00 Come out of Controlled Access, and MCC prepares to give us "tune beam"...immediately note that HRS-L Quads are all tripped off; Mark Stevens goes back down to manually reset them.

    14:30 Ready to go again. Quads reset.

    By end of shift ... still waiting for beam.